OpenStudy (imalexlowe):

Which of the following us the equation of a line perpendicular to the line y=-3/2x+4 passing through the point (3,9) A. 2x-3y=21 B. -2x+3y=21 C. 2x+3y=21 D. -2x-3y=-21

OpenStudy (dumbcow):

equation of a line \[y - y_1 = m(x-x_1)\] given point (3,9) \[y-9 = m(x-3)\] where m is slope perpendicular to -3/2 ----> take opposite reciprocal \[y-9 = \frac{2}{3} (x-3)\] \[y = \frac{2}{3} x +7\] \[3y = 2x + 21\]

OpenStudy (imalexlowe):

So A?

OpenStudy (imalexlowe):

It was B